Blossom Legler requires an estimate of the cost of goods lost by fire on March 9. Merchandise on hand on January 1 was $47,880. Purchases since January 1 were $90,720; freight-in, $4,284; purchase returns and allowances, $3,024. Sales are made at 33 1/3% above cost and totaled $150,000 to March 9. Goods costing $13,734 were left undamaged by the fire; remaining goods were destroyed.

A) Compute the cost of goods destroyed. (Round gross profit percentage and final answer to 0 decimal places, e.g. 15% or 125.)

Cost of goods destroyed

B) Compute the cost of goods destroyed, assuming that the gross profit is 33 1/3% of sales. (Round ratios for computational purposes to 5 decimal places, e.g. 78.72345% and final answer to 0 decimal places, e.g. 28,987.)

Cost of goods destroyed
Answer #1

Beginning inventory = $47,880

Purchases = $90,720

Freight in = $4,284

Purchase returns and allowances = $3,024.

Sales = $150,000

Goods costing $13,734 were left undamaged by the fire.

A)

Sales are made at 33 1/3% above cost

It means, gross profit is 33 1/3% on cost

Cost of goods sold + Gross profit = Sales

Let Cost of goods sold be $K

Gross profit = K x 1/3

= K/3

K + K/3 = 150,000

4K = 450,000

K = $112,500

Cost of goods sold = $112,500

Cost of goods destroyed = Beginning inventory + Purchases + Freight in - Purchase returns and allowances - Cost of goods sold - Goods left undamaged by the fire

= 47,880 + 90,720 + 4,284 - 112,500 - 13,734

= $16,650

B)

Gross profit is 33 1/3% of sales

Gross profit = 150,000 x 1/3

= $50,000

Cost of goods sold = Sales - Gross profit

= 150,000 - 50,000

= $100,000

Cost of goods destroyed = Beginning inventory + Purchases + Freight in - Purchase returns and allowances - Cost of goods sold - Goods left undamaged by the fire

= 47,880 + 90,720 + 4,284 - 100,000 - 13,734

= $29,150

Kindly give a positive rating if you are satisfied with the answer. Feel free to ask if you have any doubt. Thanks.
Answer #2

The answer is almost correct. You forgot to subtract the Purchase returns and allowances = $3,024. in both answers. 

source: just followed calculated it
answered by: Angy Chesler